OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. x + y = 3 2x + 2y = 10

OpenStudy (anonymous):

x + y = 3 ---(I) 2x + 2y = 10 ---(II) Rearranging (I) we get x = 3-y Now substitute x = 3-y in (II) and solve, you will get value of y put that in (I) and you will get the value of x then write the solution as an ordered pair.

OpenStudy (anonymous):

ok thankss